# Choosing the Better Bandit Algorithm under Data Sharing: When Do A/B Experiments Work?

### v1

@@ 2.3 Comparing two algorithms by their regrets
$\widehat{\text{GTE}}=R_{T}(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2})-R_{T}(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2}).$
@@
$\mathbb{E}[\widehat{\text{GTE}}]=\mathbb{E}[R_{T}(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2})]-\mathbb{E}[R_{T}(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2})].$
@@
3 Theoretical results
@@ 3.1 Preliminaries on classic bandit algorithms
[…]
@@
For both the $\epsilon$-greedy algorithm and the UCB algorithm, a higher value of $\alpha$ corresponds to more exploration, as shown by the higher exploration probability in the $\epsilon$-greedy algorithm, and a larger confidence bound in the UCB algorithm. Setting $\alpha=0$ yields the most common $\epsilon$-greedy algorithm with $\epsilon=\Theta(\frac{1}{t})$ and UCB algorithm with a confidence interval of length $\Theta(\sqrt{\log t/\lvert\mathcal{H}_{k,t}\rvert})$, both of which incur an optimal $O(\log T)$ regret. Hence, $\alpha=0$ corresponds to the optimal level of exploration. Different values for the parameter $\alpha$ capture algorithms with different regret performances, formalized in the following result.
@@
[…]
@@
The proof of this proposition is outlined in Appendix C.2 for completeness. For the $\epsilon$-greedy algorithm, the upper bound of $O(\log T)$ when $\alpha=0$ is proved in [5, Theorem 3]. The proof of the upper bound $O(T^{\alpha})$ when $0<\alpha\leq 1$ follows the same steps. For the upper bound of the UCB algorithm, our proof is similar to [21, Appendix C] which shows $O(T^{\alpha})$ regret for the non-anytime UCB algorithm using the confidence interval $\sqrt{\frac{2(T^{\alpha}-1)}{\alpha\left\lvert\mathcal{H}_{k,t}\right\rvert}}$ instead of $\sqrt{\frac{2(\lvert\mathcal{H}_{t}\rvert^{\alpha}-1)}{\alpha\left\lvert\mathcal{H}_{k,t}\right\rvert}}$.
@@ 3.2 Sign violation
[…]
@@
The proof of this theorem is provided in Section C.3. For the greedy algorithm, its regret decreases from being linear to a constant when running jointly with the $\epsilon$-greedy algorithm. For the $\epsilon$-greedy algorithm, combining Theorem 1 and Proposition 1, we find that its regret remains $\Omega(T^{\alpha})$ (or $\Omega(\log T)$ when $\alpha=0$), with or without data sharing. Thus, the expected GTE estimate $\mathbb{E}[\widehat{\text{GTE}}]$ under data sharing is $-\Omega(T^{\alpha})$ (or $-\Omega(\log T)$ when $\alpha=0$), while the true GTE is $\Theta(T)$, giving a sign reversal.
@@
The proof of this theorem is provided in Section C.4. Similar to Theorem 1, we find that the linear regret of the greedy algorithm decreases to a constant when running jointly with the UCB algorithm, while the regret of the UCB algorithm remains the same order, with or without data sharing. The expected GTE estimate under data sharing is again $-\Omega(T^{\alpha})$ (or $-\Omega(\log T)$ when $\alpha=0$) whereas the true GTE is $\Theta(T)$, indicating opposite signs between the expected GTE estimate and the true GTE.
@@
The proof of this theorem is provided in Appendix C.5. The high-probability bound (5) is a mild condition, as it concerns only the number of times that the optimal arm is pulled. For any algorithm with a sublinear regret, the optimal arm is expected to be pulled $T-o(T)$ times. The bound (5) merely requires the optimal arm to be pulled $\Omega(\log T)$ times with high probability. Theorems 1 and 2 are established by proving condition (5) for the $\epsilon$-greedy and UCB algorithms, respectively. In Appendix A, we consider a variant setting with one-way data sharing, where one algorithm runs individually using its own data, whereas the other algorithm has access to the data by both algorithms. Under this one-way data sharing, we provide additional results for the EXP3 algorithm.
@@ 3.3 Sign preservation
The previous results illustrate scenarios in which data sharing flips the sign of the expected GTE estimate. A natural follow-up question is whether there exist settings where the interference introduced by data sharing does not affect the resulting comparison, or equivalently, the sign of the expected GTE estimate. We show that if both algorithms belong to the parameterized class of $\epsilon$-greedy or UCB algorithms, then the sign of the expected GTE estimate is preserved.
@@
$\mathbb{E}\left[R_{T}(\texttt{$\epsilon$-grd}_{\alpha_{1},C}\mid(\texttt{$\epsilon$-grd}_{\alpha_{1},C},\texttt{$\epsilon$-grd}_{\alpha_{2},C}))\right]$ =$=\begin{cases}\Theta(\log T)&\text{if }\alpha_{1}=0,\\ \Theta(T^{\alpha_{1}})&\text{if }0<\alpha_{1}\leq 1,\end{cases}$
$\mathbb{E}\left[R_{T}(\texttt{$\epsilon$-grd}_{\alpha_{2},C}\mid(\texttt{$\epsilon$-grd}_{\alpha_{2},C},\texttt{$\epsilon$-grd}_{\alpha_{1},C}))\right]$ $=\Theta(T^{\alpha_{2}}).$
@@
[…]
@@
When we jointly run two UCB algorithms $\texttt{UCB}_{\alpha_{1}}$ and $\texttt{UCB}_{\alpha_{2}}$ with $10\leq\alpha_{1}<\alpha_{2}<$ 1, the expected regrets of the two algorithms under data sharing satisfy
$\mathbb{E}\left[R_{T}(\texttt{UCB}_{\alpha_{1}}\mid(\texttt{UCB}_{\alpha_{1}},\texttt{UCB}_{\alpha_{2}}))\right]$ =$=\begin{cases}O(\log T)&\text{ if }\alpha_{1}=0,\\ O(T^{\alpha_{1}})&\text{ if }0<\alpha_{1}\leq 1,\end{cases}$
$\mathbb{E}\left[R_{T}(\texttt{UCB}_{\alpha_{2}}\mid(\texttt{UCB}_{\alpha_{1}},\texttt{UCB}_{\alpha_{2}}))\right]$ $=\Theta(T^{\alpha_{2}}).$
The proof of this theorem is provided in Section C.7. The sign of the expected GTE estimate is again preserved when comparing two UCB algorithms: the one with the higher value of $\alpha$ incurs a higher regret, both with and without data sharing. Similar to the case of $\epsilon$-greedy algorithms, the reason that $\texttt{UCB}_{\alpha_{2}}$ performs worse than $\texttt{UCB}_{\alpha_{1}}$ is due to over-exploration: even under data sharing, $\texttt{UCB}_{\alpha_{2}}$ requires approximately $\Theta(T^{\alpha_{2}})$ samples of a suboptimal arm to achieve a sufficiently small confidence bound, while $\texttt{UCB}_{\alpha_{1}}$ can only contribute $O(T^{\alpha_{1}})$ such samples.
These two results together show that if both algorithms sufficiently explore to achieve sublinear regrets when running individually, the sign of the expected GTE estimate is preserved under data sharing as their performance is determined by the level of exploration, which is inherent to the algorithms and thus unaffected by sharing data. In practice, if a company carefully designs algorithms that perform reasonably well when running individually, the A/B experiments lead to the correct conclusions despite data sharing.

### v2

@@ 2.3 Comparing two algorithms by their regrets
$\widehat{\text{GTE}}=R_{T}(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))-R_{T}(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2})).$
@@
$\mathbb{E}[\widehat{\text{GTE}}]=\mathbb{E}[R_{T}(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))]-\mathbb{E}[R_{T}(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))].$
@@
3 Two regimes: sign violation and sign preservation
@@ 3.1 Preliminaries on classic bandit algorithms
    The $\epsilon$-greedy algorithm (denoted by $\epsilon$-grd) maintains a sequence of probabilities $\{\epsilon_{t}\}_{t\in[T]}$ with each $\epsilon_{t}\in[0,1]$. At each timestep $t\in[T]$, the $\epsilon$-greedy algorithm selects one of the $K$ arms uniformly at random with probability $\epsilon_{t}$, and otherwise follows the greedy algorithm with probability $(1-\epsilon_{t})$. We define a class of $\epsilon$-greedy algorithms (denoted by $\texttt{$\epsilon$-grd}_{\alpha,C}$) parameterized by a parameter $\alpha\in[0,1]$ and a constant $C>0$, where the exploration probability is defined by $\epsilon_{t}=\min\left\{1,\frac{C}{\left\lvert\mathcal{H}_{t}\right\rvert^{1-\alpha}}\right\}$.
@@
[…]
@@
The expected regret of the $\epsilon$-greedy algorithm $\texttt{$\epsilon$-grd}_{\alpha,C}$ with $C\geq\max\left\{120K,\frac{16K}{\Delta_{\textrm{min}}^{2}}\right\}$ satisfies
@@
The proof of this proposition is outlined in Appendix C.2 for completeness. For the $\epsilon$-greedy algorithm, the upper bound of $O(\log T)$ when $\alpha=0$ is proved in [5, Theorem 3]. The proof of the upper bound $O(T^{\alpha})$ when $0<\alpha\leq 1$ follows the same steps. For the upper bound of the UCB algorithm, our proof is similar to [22, Appendix C] which shows $O(T^{\alpha})$ regret for the non-anytime UCB algorithm using the confidence interval $\sqrt{\frac{2(T^{\alpha}-1)}{\alpha\left\lvert\mathcal{H}_{k,t}\right\rvert}}$ instead of $\sqrt{\frac{2(\lvert\mathcal{H}_{t}\rvert^{\alpha}-1)}{\alpha\left\lvert\mathcal{H}_{k,t}\right\rvert}}$.
3.2 A general observation
We make the following observation that connects the exploration-exploitation tradeoff of the two algorithms and the regimes of sign preservation and sign violation.
Observation 1.
When comparing two bandit algorithms, whether $\mathbb{E}[\widehat{\text{GTE}}$] and GTE have the same sign is driven by the exploration-exploitation tradeoff of the two algorithms.
  • If the worse algorithm (when run individually without data sharing) underperforms because it over-explores, then sign preservation occurs under data sharing. That is, $\mathbb{E}[\widehat{\text{GTE}}$] and GTE have the same sign.
  • If the worse algorithm (when run individually without data sharing) underperforms because it over-exploits, then sign violation occurs under data sharing. That is, $\mathbb{E}[\widehat{\text{GTE}}$] and GTE have opposite signs.
In what follows, we formally establish this result by analyzing the greedy algorithm as an example of over-exploitation, and the parameterized classes of the greedy and UCB algorithms (with $0<\alpha\leq 1$) as examples of over-exploration.
@@ 3.3 Sign violation
When we jointly run the greedy algorithm grd and the $\epsilon$-greedy algorithm $\texttt{$\epsilon$-grd}_{\alpha,C}$ with $C\geq\max\left\{80K,\frac{32K}{\Delta_{\textrm{min}}^{2}}\right\}$, the expected regrets of the two algorithms under data sharing satisfy
@@
The proof of this theorem is provided in Section C.3. For the greedy algorithm, its regret decreases from being linear to a constant when running jointly with the $\epsilon$-greedy algorithm. For the $\epsilon$-greedy algorithm, combining Theorem 1 and Proposition 1, we find that its regret remains $\Omega(T^{\alpha})$ (or $\Omega(\log T)$ when $\alpha=0$), with or without data sharing. Thus, the expected GTE estimate $\mathbb{E}[\widehat{\text{GTE}}$] under data sharing is $\Omega(T^{\alpha})$ (or $\Omega(\log T)$ when $\alpha=0$), while the true GTE is $\Theta(T)$, giving a sign reversal.
@@
The proof of this theorem is provided in Section C.4. Similar to Theorem 1, we find that the linear regret of the greedy algorithm decreases to a constant when running jointly with the UCB algorithm, while the regret of the UCB algorithm remains the same order, with or without data sharing. The expected GTE estimate under data sharing is again $\Omega(T^{\alpha})$ (or $\Omega(\log T)$ when $\alpha=0$) whereas the true GTE is $\Theta(T)$, indicating opposite signs between the expected GTE estimate and the true GTE.
@@
The proof of this theorem is provided in Appendix C.5. The high-probability bound (5) is a mild condition, as it concerns only the number of times that the optimal arm is pulled. For any algorithm with a sublinear regret, the optimal arm is expected to be pulled To(T)T-o(T) times. The bound (5) merely requires the optimal arm to be pulled $\Omega(\log T)$ times with high probability. Theorems 1 and 2 are established by proving condition (5) for the $\epsilon$-greedy and UCB algorithms, respectively. In Appendix A, we consider a variant setting with one-way data sharing, where one algorithm runs individually using its own data, whereas the other algorithm has access to the data by both algorithms. Under this one-way data sharing, we provide additional results for the EXP3 algorithm.
@@ 3.4 Sign preservation
The previous results illustrate scenarios in which data sharing flips the sign of the expected GTE estimate. A natural follow-up question is whether there exist settings where data sharing does not affect the resulting comparison. We show that if both algorithms belong to the parameterized class of $\epsilon$-greedy or UCB algorithms, then the sign of the expected GTE estimate is preserved.
@@
When we jointly run two algorithms $\mathcal{A}_{\alpha_{1}}\in\{\texttt{$\epsilon$-grd}_{\alpha_{1},C},\,\texttt{UCB}_{\alpha_{1}}\}$, $\mathcal{A}_{\alpha_{2}}\in\{\texttt{$\epsilon$-grd}_{\alpha_{2},C},\,\texttt{UCB}_{\alpha_{2}}\}$, with $0\leq\alpha_{1}<\alpha_{2}\leq 1$ and $C\geq\max\{120K,\,32K/\Delta_{\textrm{min}}^{2}\}$, the expected regrets of the two algorithms under data sharing satisfy
$\mathbb{E}\left[R_{T}(\mathcal{A}_{\alpha_{1}}\mid(\mathcal{A}_{\alpha_{1}},\mathcal{A}_{\alpha_{2}}))\right]=\begin{cases}O(\log T),&\text{if }\alpha_{1}=0,\\ O(T^{\alpha_{1}}),&\text{if }0<\alpha_{1}\leq 1,\end{cases}$
$\mathbb{E}\left[R_{T}(\mathcal{A}_{\alpha_{2}}\mid(\mathcal{A}_{\alpha_{1}},\mathcal{A}_{\alpha_{2}}))\right]=\Theta(T^{\alpha_{2}}).$
The proof of this theorem is provided in Section C.6. Looking at Theorem 4 together with Proposition 1, we find that under data sharing, the regret of $\mathcal{A}_{\alpha_{1}}$ remains $O(T^{\alpha_{1}})$ (or $O(\log T)$ when $\alpha_{1}=0$), while the regret of $\mathcal{A}_{\alpha_{2}}$ remains $\Theta(T^{\alpha_{2}})$. As a result, the sign of the expected GTE estimate is preserved when comparing two such algorithms: the one with the larger value of $\alpha$ incurs the higher regret, both with and without data sharing. Intuitively, $\mathcal{A}_{\alpha_{2}}$ performs worse because it over-explores: it explores more than $\mathcal{A}_{\alpha_{1}}$ and, in particular, more than the optimal rate, regardless of data sharing. This excess exploration accounts for the majority of its regret. For $\epsilon$-greedy, the probability of selecting each arm under $\mathcal{A}_{\alpha_{2}}$ is always bounded below by its exploration rate $\epsilon_{t}$, so $\texttt{$\epsilon$-grd}_{\alpha_{2},C}$ continues to over-explore under data sharing. For UCB, $\texttt{UCB}_{\alpha_{2}}$ requires approximately $\Theta(T^{\alpha_{2}})$ samples of a suboptimal arm to achieve a sufficiently small confidence bound, whereas the other algorithm $\mathcal{A}_{\alpha_{1}}$ contributes only $O(T^{\alpha_{1}})$ such samples; thus, it also over-explores under data sharing.
The result shows that if both algorithms sufficiently explore to achieve sublinear regrets when running individually, the sign of the expected GTE estimate is preserved under data sharing as their performance is determined by the level of exploration, which is inherent to the algorithms and thus unaffected by sharing data.
4 Detecting sign violation with ramp-up experiments
In practice, we often do not know a priori whether an algorithm over-explores or over-exploits before running an experiment (which is why we need experiments to start with!). That is, we cannot determine whether an experiment lies in the regime of sign preservation or sign violation, and thus do not know how to interpret the comparison results from the experiment. In this section, we provide a procedure to determine whether an experiment lies in the sign violation or sign preservation regime.
4.1 Detection procedure
We propose to detect potential sign violation using ramp-up A/B experiments. We first extend the previous data sharing scheme with a fixed 50–50 split in Algorithm 1 to a general Bernoulli assignment in Algorithm 2, where each user is independently assigned to treatment with probability $p$. A ramp-up experiment then consists of a sequence of tests with treatment probabilities $0<p_{1}<p_{2}<\cdots$, where Algorithm 2 is first run with treatment probability $p_{1}$, and the treatment probability is subsequently increased to $p_{2}$, and then $p_{3}$, and so on. In online experimentation, A/B experiments are often run in this ramp-up fashion, starting with a low treatment probability, typically for a new algorithm that has not been deployed before, in order to monitor performance and detect unintended side effects at low exposure levels before broader deployment [41].
1
3 for t=1, …, 2T do
4   Sample $W_{t}\sim\text{Bern}(p)$
5 if $W_{t}=0$ then
6      Algorithm $\mathcal{A}_{1}$ selects arm $I=\mathcal{A}_{1}(\mathcal{H}_{t})$ and observes reward $Y_{t}\sim D_{I}$
7 else
8     Algorithm $\mathcal{A}_{2}$ selects arm $I=\mathcal{A}_{2}(\mathcal{H}_{t})$ and observes reward $Y_{t}\sim D_{I}$
9   end if
10  Add the reward sample to the history:
$\mathcal{H}_{I,t}\leftarrow\mathcal{H}_{I,t}\cup\{Y_{t}\}$
$\mathcal{H}_{t+1}\leftarrow\mathcal{H}_{t}$
11 end for
Algorithm 2 Running two algorithms $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ under data sharing, with a Bernoulli assignment (probability (1p)(1-p) for $\mathcal{A}_{1}$ and probability $p$ for $\mathcal{A}_{2}$).
We now propose a simple warning system that alerts experimenters when they may be operating in a sign violation regime. The key idea is to monitor how the overall average outcome evolves as the treatment probability increases during a ramp-up experiment. Concretely, suppose that at a baseline allocation, Algorithm 2 appears to outperform Algorithm 1. As the share of traffic assigned to Algorithm 2 increases, one would typically expect the overall average outcome to improve. A decline in the overall average outcome as the traffic share of Algorithm 2 increases suggests a potential sign violation.
Formally, consider a Bernoulli assignment with treatment probability $p$. At each timestep $t$, let $W_{t}\sim\text{Bern}(p)$ denote the treatment assignment, where $W_{t}=1$ corresponds to Algorithm 2 and $W_{t}=0$ to Algorithm 1. Define the inverse probability weighted (IPW) estimators of the mean outcome for each algorithm as
$\widehat{Y}^{p}(\mathcal{A}_{2})=\frac{\sum_{t\in[2T]}Y_{t}W_{t}}{2Tp},\qquad\widehat{Y}^{p}(\mathcal{A}_{1})=\frac{\sum_{t\in[2T]}Y_{t}(1-W_{t})}{2T(1-p)},$
and the overall average outcome as
$\overline{Y}^{p}=\frac{1}{2T}\sum_{t\in[2T]}Y_{t}.$
The detection rule is as follows:
  • In a ramp-up experiment, if increasing the traffic assigned to the empirically better algorithm results in a decline in the overall average outcome, we flag this behavior as a potential sign violation and recommend further investigation.
Mathematically, we flag a potential sign violation if there exist treatment probabilities $p<p^{\prime}$ such that
$\operatorname*{sign}\left(\widehat{Y}^{0.5}(\mathcal{A}_{2})-\widehat{Y}^{0.5}(\mathcal{A}_{1})\right)\neq\operatorname*{sign}\left(\overline{Y}^{p^{\prime}}-\overline{Y}^{p}\right),$ (6)
The intuition behind this diagnostic is as follows. When increasing the probability on the better algorithm worsens the average outcome, a likely cause is that the better algorithm free rides the other algorithm, and once the traffic of the other algorithm decreases, the level of exploration is no longer sufficient for the (seemingly) better algorithm to perform well. For example, consider the UCB algorithm with the greedy algorithm. Under data sharing, the greedy algorithm appears to perform better than the UCB algorithm, yielding a higher average outcome. However, when the greedy algorithm is assigned a large share of traffic, the UCB algorithm now only takes a very small share of traffic and no longer provides sufficient exploration that is needed by the greedy algorithm, resulting in a high regret and, consequently, a low average outcome.
In contrast, consider a second example comparing two UCB algorithms, $\texttt{UCB}_{0}$ and $\texttt{UCB}_{0.5}$. $\texttt{UCB}_{0}$ already explores sufficiently on its own for any treatment probability. Increasing the share of $\texttt{UCB}_{0}$ therefore reduces unnecessary exploration by $\texttt{UCB}_{0.5}$, which reduces the total regret and improves the average outcome. The sign terms on the left- and right-hand sides of the detection criterion (6) coincide, indicating sign preservation.
4.2 Theoretical justifications of the detection procedure
We now provide theoretical justifications for the proposed detection procedure. In Algorithm 2 under a Bernoulli assignment with probability $p$, we define the regrets of the two algorithms:
$R^{p}(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))$ $=\sum_{t\in[2T]:W_{t}=0}(\mu^{*}-Y_{t}),$
$R^{p}(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))$ $=\sum_{t\in[2T]:W_{t}=1}(\mu^{*}-Y_{t}),$
with $W_{t}\sim\text{Bern}(p)$. We further define the total regret:
$R^{p}(\texttt{tot}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))$ $=R^{p}(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))+R^{p}(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))$
$=2T\mu^{*}-\sum_{t\in[2T]}Y_{t}.$
We note that the detection procedure (6) is equivalent to the following criterion. We flag a potential sign violation if there exist $p<p^{\prime}$ such that
$\operatorname*{sign}\big(R^{0.5}(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))-R^{0.5}(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2})\big)\neq\operatorname*{sign}\big(R^{p^{\prime}}(\texttt{tot}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))-R^{p}(\texttt{tot}\mid(\mathcal{A}_{1},\mathcal{A}_{2})\big).$ (7)
In what follows, we build upon our results from Section 3, and use the $\epsilon$-greedy and UCB algorithms as examples to show that (1) a sign violation caused by running the greedy algorithm jointly with the $\epsilon$-greedy or UCB algorithm can be correctly detected by (7); (2) under a sign preservation by running a pair of $\epsilon$-greedy or UCB algorithms, no warning is issued according to (7).
First, we consider the regime of sign violation. According to (7), any non-monotonic behavior in the total regret as the treatment probability increases triggers a warning. We now show that when the greedy algorithm runs jointly with the $\epsilon$-greedy or UCB algorithm, this warning is indeed triggered. Suppose Algorithm 1 is either the $\epsilon$-greedy or UCB algorithm, while Algorithm 2 is the greedy algorithm. The results in Theorems 1 and 2 can be readily extended to a Bernoulli assignment with probability $0.5$ to show that Algorithm 2 performs better under data sharing. Now think about a ramp-up procedure with $p$ increasing from 0 to $1$. When $p=0$, the system only runs Algorithm 1, incurring a low regret. When $p=1$, the system reduces purely to the greedy algorithm, which incurs a linear regret. Therefore, the total regret must increase at some point. This corresponds precisely to the situation in which increasing the treatment probability of the seemingly better performing algorithm leads to an increase in the total regret, thereby triggering a warning for a potential sign violation.
Now we move on to the regime of sign preservation. We show that when running a pair of $\epsilon$-greedy or UCB algorithms with parameters $0\leq\alpha_{1}<\alpha_{2}\leq 1$, the regret is monotonically increasing in the probability of the worse algorithm. Specifically, we consider treatment probabilities of the form
$p=T^{\beta-1},$
for varying values of $\beta\in(0,1)$. The following two theorems provide the regret rates in terms of $\beta$ for the $\epsilon$-greedy and UCB algorithms, respectively. We start with the following results for $\epsilon$-greedy algorithms.
Theorem 5.
Consider running two $\epsilon$-greedy algorithms, $\texttt{$\epsilon$-grd}_{\alpha_{1},C}$ and $\texttt{$\epsilon$-grd}_{\alpha_{2},C}$ with $0\leq\alpha_{1}<\alpha_{2}\leq 1$ and $C\geq\max\{120K,\;32K/\Delta_{\textrm{min}}^{2}\}$. Suppose the traffic allocation percentage to $\texttt{$\epsilon$-grd}_{\alpha_{2},C}$ is $p=T^{\beta-1}$ for some $\beta\in(0,1)$. Then the expected total regret under data sharing satisfies
$\mathbb{E}\left[R^{p}(\texttt{tot}\mid(\texttt{$\epsilon$-grd}_{\alpha_{1},C},\texttt{$\epsilon$-grd}_{\alpha_{2},C})\right]=\begin{cases}\Theta(\log T)&\text{if }\alpha_{1}=0\text{ and }\alpha_{2}+\beta\leq 1,\\ \Theta(T^{\max\left\{\alpha_{1},\alpha_{2}+\beta-1\right\}})&\text{otherwise.}\end{cases}$ (8)
The proof of this theorem is provided in Section C.7.1. Now we provide the results for UCB algorithms.
@@ 4.2 Theoretical justifications of the detection procedure
((a))
@@
((b))
Figure 3: Total regret as a function of the traffic allocation parameter $\beta$.
@@
Consider running two UCB algorithms, $\texttt{UCB}_{\alpha_{1}}$ and $\texttt{UCB}_{\alpha_{2}}$, with parameters $0\leq\alpha_{1}<\alpha_{2}\leq 1$. Suppose the traffic allocation percentage to $\texttt{UCB}_{\alpha_{2}}$ is $p=T^{\beta-1}$ for some $\beta\in(0,1)$. Then the expected total regret under data sharing satisfies
$\mathbb{E}\left[R^{p}(\texttt{tot}\mid(\texttt{UCB}_{\alpha_{1}},\texttt{UCB}_{\alpha_{2}},p)\right]=\Theta(T^{\max\left\{\alpha_{1},\min\{\alpha_{2},\beta\}\right\}}).$ (9)
The proof of this theorem is provided in Section C.7.2. Taken together, these two theorems show that the total regret is monotonically increasing as the probability of the worse algorithm increases, as illustrated in Figure 3. Specifically, the algorithm with exploration parameter $\alpha_{1}$ outperforms the algorithm with parameter $\alpha_{2}$. With increasing treatment probability (i.e., increasing $\beta$) for the worse algorithm (i.e., $\texttt{$\epsilon$-grd}_{\alpha_{2},C}$ or $\texttt{UCB}_{\alpha_{2},C}$), the total regret increases.
As an aside, while the $\epsilon$-greedy and UCB algorithms behave similarly in our previous results under a fixed 50-50 split, their regret rates become different in a ramp-up experiment as seen from (8) and (9), due to the following reason. The exploration level of the $\epsilon$-greedy algorithm is determined by its parameter $\alpha$, whereas the exploration level of the UCB algorithm can also be affected by data sharing: If the other algorithm exploits and pulls the optimal arm excessively, the confidence interval of the optimal arm becomes very small, pushing the UCB algorithm to explore even more. Thus, the pair of UCB algorithms reach a high regret of order $T^{\alpha_{2}}$ faster, before $\texttt{UCB}_{\alpha_{2}}$ is allocated the full traffic.